OpenStudy (anonymous):

find the x intercepts and the y intercepts to sketch a graph. x +y = 3

OpenStudy (anonymous):

METHOD 1 to find y-intercept put x=0 and to find x-intercpets put y=0 METHOD 2 reduce the equation to the intercept form i.e (x/a)+(y/b)=1 thn x-intercept is a unit and y-intercept is b unit

OpenStudy (anonymous):

here x +y = 3 can be reduced to (x/3) +(y/3)=1 thus x-intercpt is 3 unit (pt on x-axis (3,0)) and y-intercept is 3 unit (pt on y-axis (0,3))

OpenStudy (anonymous):

can you show me using method 1... i am confused thanks.